A random variable x has the cumulative distribution function given as



f(x) = 8>x<7



0; for x<1



X^2 -2x+2



2; for 1≤x2



1; for x≥2



Calculate the variance of x.




1
Expert's answer
2021-12-01T17:49:39-0500
"F(x)= \begin{cases}
 0 &x<1 \\
 x^2-2x+1 &1\leq x<2 \\
 1 & x\geq2
\end{cases}"

"f(x)= \begin{cases}
 2x-2 &1\leq x\leq2 \\
 0 & otherwise
\end{cases}"

Check


"\displaystyle\int_{-\infin}^{\infin}f(x)dx=\displaystyle\int_{1}^{2}(2x-2)dx"

"=[x^2-2x]\begin{matrix}
 2 \\
 1
\end{matrix}=4-4-(1-2)=1"

"E(X)=\displaystyle\int_{-\infin}^{\infin}xf(x)dx=\displaystyle\int_{1}^{2}x(2x-2)dx"

"=[\dfrac{2x^3}{3}-x^2]\begin{matrix}
 2 \\
 1
\end{matrix}=\dfrac{16}{3}-4-(\dfrac{2}{3}-1)=\dfrac{5}{3}"

"E(X^2)=\displaystyle\int_{-\infin}^{\infin}x^2f(x)dx=\displaystyle\int_{1}^{2}x^2(2x-2)dx"

"=[\dfrac{x^4}{2}-\dfrac{2x^3}{3}]\begin{matrix}
 2 \\
 1
\end{matrix}=8-\dfrac{16}{3}-(\dfrac{1}{2}-\dfrac{2}{3})=\dfrac{17}{6}"

"Var(X)=E(X^2)-[E(X)]^2"

"=\dfrac{17}{6}-[\dfrac{5}{3}]^2=\dfrac{1}{18}"